# PromptReps: Prompting Large Language Models to Generate Dense and Sparse Representations for Zero-Shot Document Retrieval

### v1

@@ Abstract
The current use of large language models (LLMs) for zero-shot document ranking follows one of two ways: 1) prompt-based re-ranking methods, which require no further training but are feasible for only re-ranking a handful of candidate documents due to the associated computational costs; and 2) unsupervised contrastive trained dense retrieval methods, which can retrieve relevant documents from the entire corpus but require a large amount of paired text data for contrastive training. In this paper, we propose PromptReps, which combines the advantages of both categories: no need for training and the ability to retrieve from the whole corpus. Our method only requires prompts to guide an LLM to generate query and document representations for effective document retrieval. Specifically, we prompt the LLMs to represent a given text using a single word, and then use the last token’s hidden states and the corresponding logits associated to the prediction of the next token to construct a hybrid document retrieval system. The retrieval system harnesses both dense text embedding and sparse bag-of-words representations given by the LLM. Our experimental evaluation on the BEIR zero-shot document retrieval datasets illustrates that this simple prompt-based LLM retrieval method can achieve a similar or higher retrieval effectiveness than state-of-the-art LLM embedding methods that are trained with large amounts of unsupervised data, especially when using a larger LLM. [^1] Code for fully reproducing the results is available at https://github.com/ielab/PromptReps.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT4 and LLaMA, which are pretrained on massive corpora and finetuned to follow user instructions, have strong zero-shot natural language understanding capabilities OpenAI (2023); Touvron et al. (2023). Via prompting, LLMs excel in various text generation tasks such as question answering, writing assistance, and conversational agent Hendrycks et al. (2021); Liu et al. (2023). Inspired by the success of LLMs on natural language understanding tasks, several studies have explored the potential of using LLMs to perform unsupervised document ranking task.
Figure 1: Overview of PromptReps, LLMs are prompted to generate dense and sparse text representation at the same time, which is then used to build search indices.
@@
In this work, we propose a new zero-shot LLM-based document retrieval method called PromptReps. We demonstrate that LLMs can be directly prompted to produce query and document embeddings, which can serve as effective text representations for neural retrieval systems. Specifically, we prompt an LLM by asking it to use a single word to represent a query or a document. Then, we extract the last layer’s hidden state of the last token in the prompt as the dense representation of the input text. Simultaneously, we utilize the logits associated with predicting the subsequent token to form a sparse representation. As illustrated in Figure 1, through a single forward pass, we generate text representations for a document that can be indexed for dense, sparse, or hybrid search architectures.
We evaluate our approach on the BEIR datasets Thakur et al. (2021) to perform zero-shot retrieval. We utilize various open-source large language models and find that the effectiveness of zero-shot retrieval increases with the scaling of LLM size. Results show that PromptReps outperforms previous LLM-based embedding methods. Of key importance is that our method is the first LLM-based method that can effectively perform full corpus retrieval while at the same time not requiring contrastive training, demonstrating that prompt engineering for generative LLMs is capable of generating robust representations for retrieval.
@@ 2.1 Supervised Neural Retrievers
Neural retrievers based on bi-encoder architecture bring significant improvement over traditional heuristic retrievers such as BM25. Dense retrievers such as DPR Karpukhin et al. (2020), ANCE Xiong et al. (2021), ColBERT Khattab and Zaharia (2020), encode text into low-dimensional dense vectors and conduct search with (approximate) nearest neighbor search. On the other hand, sparse neural retrievers such as DeepImpact Mallia et al. (2021), uniCOIL Lin and Ma (2021), TILDE Zhuang and Zuccon (2021b, a), and SPLADE Formal et al. (2021), encode text into high-dimensional sparse vectors as bag-of-words representations and conduct search in an inverted index. Recent work has also explored fine-tuning LLMs as dense retrievers such as RepLLaMA Ma et al. (2023a) and LLaRA Liao et al. (2024). A hybrid neural retrieval system refers to a system that combines the rankings provided by both dense and sparse retrievers, often resulting in an enhanced final ranking Lin and Ma (2021). All these retrievers are trained with supervised relevance judgment data (e.g., MS MARCO Bajaj et al. (2018)) using contrastive learning. Our work instead focuses on building a hybrid neural retrieval system with zero-shot dense and sparse document representations without supervised contrastive learning. This is particularly valuable for applications where no relevance judgements are available — often the case in real work retrieval settings.
@@ 2.2 Unsupervised Neural Retrievers
There have also been attempts on training effective neural retrievers without relying on human relevance judgments. Methods such as Contriever Izacard et al. (2022) and E5 Wang et al. (2024a), train dense retriever with large-scale pseudo query-document pairs to build unsupervised (synthetic) training data. LLMs have also been adapted as unsupervised text embedding models for first-stage document retrieval. For instance, HyDE Gao et al. (2023a) enhances query representations for an unsupervised retriever by replacing the original query with LLM-generated hypothetical documents. More recent work has focused on directly converting generative LLMs into a text-embedding models with unsupervised contrastive pre-training. These works, such as E5-Mistral-Inst Wang et al. (2024b) and Gecko Lee et al. (2024), use large-scale weakly supervised paired text data or LLM-generated query-document pair data to perform contrastive training on top of LLMs. LLM2Vec BehnamGhader et al. (2024), on the other hand, conducts further masked next token prediction pre-training with bidirectional attention, and SimCSE Gao et al. (2021) trains on raw text data to transform LLMs into text encoders. Although no labeled data is used, these methods require synthetic or unsupervised paired text data to perform contrastive pre-training. Our method instead relies solely on prompt engineering to transform LLM into a robust text encoder for document retrieval without any extra training.
2.3 Prompting LLMs for document re-ranking
Inspired by the prompt-following capacity of LLMs, recent works have explored prompting LLMs for document reranking. For instance, UPR Sachan et al. (2022) ranks documents pointwise by prompting the LLM to generate a relevant query for a given document and rank documents based on the likelihood of generating the query. RankGPT Sun et al. (2023) and LRL Ma et al. (2023b) propose to re-rank a list of documents at once and generate permutations for the reordered list. Pairwise Qin et al. (2024) and Setwise Zhuang et al. (2023b) prompting methods have also been explored to improve effectiveness and efficiency in the LLM re-ranking pipeline. These methods are only feasible for re-ranking a handful of candidate documents, thus limited to second-stage document re-ranking. In contrast, our approach utilizes prompts to construct the first-stage retrievers.
@@ 2.4 Prompting LLM for sentence embeddings
The methods most similar to ours are work on prompting LLMs to generate sentence embeddings for semantic textual similarity (STS) tasks Jiang et al. (2023b); Lei et al. (2024); Zhang et al. (2024). These works developed an Explicit One-word Limitation (EOL) prompt, which also instructs LLMs to represent a sentence with one word. However, these works only evaluate such prompts on STS datasets, and their effectiveness on information retrieval datasets with large document corpora is unknown. Additionally, these methods only represent texts with dense embeddings from the hidden states; our method instead generates sparse text representations simultaneously to build a hybrid retrieval system. In fact, our results show that dense embeddings alone perform poorly for document retrieval tasks with some LLMs, but sparse representations are much more robust, and the best retrieval effectiveness is achieved with the hybrid retrieval system with scaled model size.
@@ 3 PromptReps
Unlike previous works that leverage LLMs for document ranking, which are limited to document re-ranking tasks with prompts or rely on contrastive learning to transform a generative LLM into an embedding model for document retrieval, in this work, we aim to directly prompt LLMs to generate both dense embedding representations and sparse bag-of-words representations for document retrieval without any form of extra training effort. To achieve this aim, we devise the following prompt as the input text for LLMs:
<System> You are an AI assistant that can understand human language. <User> Passage "[text]". Use one most important word to represent the passage in retrieval task. Make sure your word is in lowercase. <Assistant> The word is: "
where <System> <User> and <Assistant> are LLM pre-defined conversational prefix tokens and [text] is the placeholder for passage text. When using this prompt for text generation, the language model needs to find a single word in its token vocabulary that can best represent the given passage to generate. However, since there could be multiple words to represent the passage, there might be multiple tokens in the vocabulary that have a high probability of being sampled by the language model. Such a distribution over the vocabulary, which often refers to "logits", could potentially provide a good representation of the given passage. In addition, since the logits are computed by the last layer hidden state [^2] Often through dot product between last hidden state with all token embeddings. of the last input token (‘"’), which is a dense vector embedding, it can also serve as a dense representation of the passage.
@@
Specifically, during the document indexing phase, we pass all the documents with our prompt into the LLM to get output hidden states and logits. To build a sparse retrieval pipeline with logits, we first need to sparsify the logits representation to be able to perform efficient sparse retrieval. This is because logits originally had values for all tokens in the vocabulary, essentially forming dense vectors with dimensions equal to the vocabulary size. To sparsify the logit representations for sparse retrieval, we perform following steps:
@@
    Finally, the logits were quantized by simply multiplying the original values by 100 and taking the integer operation on that, and the obtained values represent the weights of corresponding tokens.
With these adjustments, the logits representations of documents are heavily sparsified, allowing for efficient sparse retrieval with a standard inverted index.
For dense retrieval, we directly use the hidden states as the embedding of the documents. For indexing these embeddings, we simply normalize all the embeddings and add them into an Approximate Nearest search (ANN) vector index.
At query time, we process the queries exactly the same as the documents, with the only exception being that the term "passage" in the prompt is replaced with "query". The dense representation of the query is utilized for semantic search via the ANN index, while the sparse representation of the query is employed for exact term matching via the inverted index. Following previous work Wang et al. (2021), we compute the final document scores by applying min-max normalization to both dense and sparse document scores. These normalized scores are then linearly interpolated with equal weights to produce the final document scores.
Table 1: nDCG@10 scores of BEIR 13 publicly available datasets. The best scores of methods without interpolating with BM25 are highlighted in bold. Scores of E5-PTlarge and LLM2Vec are taken directly from the original papers.
Contrastive pre-training PromptReps(ours)
LLM - BERT-330M Mistral-7B Mistral-7B-Instruct Phi3-mini-3.8B-Instruct Llama3-8B-Instruct Llama3-70B-Instruct
Dataset BM25 E5-PTlarge LLM2Vec Dense Sparse Hybrid Dense Sparse Hybrid Dense Sparse Hybrid Dense Sparse Hybrid +BM25
arguana 39.70 44.4 51.00 21.02 32.41 32.81 38.00 21.51 35.68 32.89 24.25 35.47 37.38 26.15 38.29 42.28
climate-fever 16.51 15.7 22.97 5.31 11.17 10.57 18.99 11.39 21.12 23.47 9.52 22.20 21.21 12.01 22.56 22.96
dbpedia 31.80 37.1 25.48 8.23 22.80 17.87 26.09 24.72 32.21 30.95 28.11 36.30 31.88 28.69 36.59 40.48
fever 65.13 68.6 45.11 20.03 50.07 47.23 44.18 51.73 64.92 59.43 51.60 71.61 57.62 54.77 71.52 76.37
fiqa 23.61 43.2 27.24 8.31 18.63 15.75 21.08 20.40 27.01 24.79 19.66 29.54 29.99 22.22 33.34 33.20
hotpotqa 63.30 52.2 54.54 2.98 32.15 20.62 25.15 42.75 46.62 28.07 45.82 51.27 30.94 44.91 50.98 66.42
nfcorpus 32.18 33.7 27.16 14.41 27.81 23.55 27.80 29.38 33.25 26.74 27.48 31.36 31.93 29.35 34.03 36.19
nq 30.55 41.7 34.16 7.12 19.93 16.62 26.47 25.57 34.78 35.02 28.34 41.62 43.38 30.76 45.98 44.22
quora 78.86 86.1 84.40 53.38 61.95 66.65 80.71 67.79 82.11 77.47 69.02 82.47 75.27 71.57 81.81 84.84
scidocs 14.90 21.8 15.35 6.54 11.42 10.28 13.35 12.24 15.58 16.77 11.52 16.93 18.31 13.22 17.96 17.73
scifact 67.89 72.3 68.67 22.71 55.27 48.15 51.69 59.94 64.86 54.26 59.06 65.61 58.92 61.53 67.41 72.09
trec-covid 59.47 61.8 55.66 34.94 62.03 53.79 51.16 56.53 62.48 63.54 55.19 68.54 71.47 63.19 77.13 77.94
touche 44.22 19.8 6.54 9.59 19.47 18.36 12.75 19.17 19.58 14.92 16.46 19.24 18.84 20.12 22.85 36.21
avg 43.70 44.61 39.86 16.51 32.70 29.40 33.65 34.09 41.55 37.48 34.31 44.01 40.55 36.80 46.19 50.07

### v2

@@ Abstract
Utilizing large language models (LLMs) for zero-shot document ranking is done in one of two ways: 1) prompt-based re-ranking methods, which require no further training but are only feasible for re-ranking a handful of candidate documents due to computational costs; and 2) unsupervised contrastive trained dense retrieval methods, which can retrieve relevant documents from the entire corpus but require a large amount of paired text data for contrastive training. In this paper, we propose PromptReps, which combines the advantages of both categories: no need for training and the ability to retrieve from the whole corpus. Our method only requires prompts to guide an LLM to generate query and document representations for effective document retrieval. Specifically, we prompt the LLMs to represent a given text using a single word, and then use the last token’s hidden states and the corresponding logits associated with the prediction of the next token to construct a hybrid document retrieval system. The retrieval system harnesses both dense text embedding and sparse bag-of-words representations given by the LLM. We further explore variations of this core idea that consider the generation of multiple words, and representations that rely on multiple embeddings and sparse distributions. Our experimental evaluation on the MSMARCO, TREC deep learning and BEIR zero-shot document retrieval datasets illustrates that this simple prompt-based LLM retrieval method can achieve a similar or higher retrieval effectiveness than state-of-the-art LLM embedding methods that are trained with large amounts of unsupervised data, especially when using a larger LLM. [^1] Code for fully reproducing the results is available at https://github.com/ielab/PromptReps.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT4 and LLaMA, which are pretrained on massive corpora and finetuned to follow user instructions, have strong zero-shot natural language understanding capabilities OpenAI (2023); Touvron et al. (2023). Via prompting, LLMs excel in various text generation tasks such as question answering, writing assistance, and conversational agent Hendrycks et al. (2021); Liu et al. (2023). Inspired by the success of LLMs on natural language understanding tasks, research has explored the potential of using LLMs to perform unsupervised document ranking.
Figure 1: Overview of PromptReps. LLMs are prompted to simultaneously generate dense and sparse representations, which are then used to build search indices.
@@
In this work, we propose a new zero-shot LLM-based document retrieval method called PromptReps. We demonstrate that LLMs can be directly prompted to produce query and document embeddings, which can serve as effective text representations for neural retrieval systems. Specifically, we prompt an LLM by asking it to use a single word to represent a query or a document. Then, we extract the last layer’s hidden state of the last token in the prompt as the dense representation of the input text. Simultaneously, we utilize the logits associated with predicting the subsequent token to form a sparse representation. As illustrated in Figure 1, through a single forward pass, we generate text representations for a document that can be indexed for dense, sparse, or hybrid search architectures. We also explore alternative representations in addition to the core idea in this paper, where we generate multiple words, and use multiple embeddings to represent an item (Figures 3 and 4).
Our empirical evaluation on multiple datasets show that PromptReps can achieve a similar or higher zero-shot retrieval effectiveness than previous trained LLM-based embedding methods, especially when a large LLM is utilized. Of key importance is that our method is the first LLM-based method that can effectively perform full corpus retrieval while at the same time not requiring contrastive training, demonstrating that prompt engineering for generative LLMs is capable of generating robust representations for retrieval.
@@ 2.1 Supervised Neural Retrievers
Neural retrievers based on the bi-encoder architecture bring significant improvements over traditional best-match retrievers such as BM25. Dense retrievers such as DPR Karpukhin et al. (2020), ANCE Xiong et al. (2021), ColBERT Khattab and Zaharia (2020), are based on encoder-only language models and encode text into low-dimensional dense vectors, conducting search with (approximate) nearest neighbor search. On the other hand, sparse neural retrievers such as DeepImpact Mallia et al. (2021), uniCOIL Lin and Ma (2021), TILDE Zhuang and Zuccon (2021c, b), and SPLADE Formal et al. (2021), also based on encoder-only language models, encode text into high-dimensional sparse vectors as bag-of-words representations, conducting search in an inverted index. Recent work has also explored fine-tuning generative LLMs as dense retrievers such as RepLLaMA Ma et al. (2023a) and LLaRA Liao et al. (2024). A hybrid neural retrieval system refers to a system that combines the rankings provided by both dense and sparse retrievers, often resulting in an enhanced final ranking Lin and Ma (2021); Wang et al. (2021a).
All these retrievers are trained with supervised relevance judgment data (e.g., MS MARCO Bajaj et al. (2018)) using contrastive learning. Our work instead focuses on building a hybrid neural retrieval system with zero-shot dense and sparse document representations without supervised contrastive learning and based on generative LLMs. This capability has two implications: (1) no contrastive training is required, which is expensive when applied to LLMs with several billions parameters, and (2) no human-labelled training data is required, which may be laborious and expensive to obtain. With regards to the first point, Wang et al. (2024b) reported that the training of E5-mistrail (7B parameters) took about 18 hours on 32 V100 GPUs, for an approximate cost of USD \$2,300222Based on 4 On-Demand p3dn.24xlarge instances, June 2024., emissions of $\approx$5.6 kgCO2e and consumption of $\approx$37.7 L of water for the associated cooling activities [^3] Emissions and water consumption estimates obtained using the frameworks of Scells et al. (2022); Zuccon et al. (2023).. Scaling this training to more and larger LLMs, and more data, will consequently further increase costs. Our proposed method does not incurr these additional contrastive pre-training costs. With regards to the second point, dense retrievers have shown to have poor generalisability when applied to data out-of-domain or out-of-task compared to the data used for contrastive training Thakur et al. (2021); Zhuang and Zuccon (2021a, 2022); Ren et al. (2023); Lin et al. (2023); Lupart et al. (2022). In presence of shift in data between training and deployment, retrieval losses can be significant: dense retrieval effectiveness can plummet far below that of best-match models like BM25 Khramtsova et al. (2023, 2024). The acquisition of in-domain/in-task training data can be costly, laborious and often impractical/impossible especially in domain-specific applications and when dealing with sensitive, private data.
@@ 2.2 Unsupervised Neural Retrievers
There have also been attempts at training effective neural retrievers without relying on human relevance judgments. Methods such as Contriever Izacard et al. (2022) and E5 Wang et al. (2024a), train a dense retriever with large-scale pseudo query-document pairs to build unsupervised (synthetic) training data. LLMs have also been adapted as unsupervised text embedding models for first-stage document retrieval. For instance, HyDE Gao et al. (2023a) enhances query representations for an unsupervised retriever by replacing the original query with LLM-generated hypothetical documents. More recent work has focused on directly converting generative LLMs into a text-embedding model with unsupervised contrastive pre-training. Methods like E5-Mistral-Inst Wang et al. (2024b) and Gecko Lee et al. (2024) use large-scale weakly supervised paired text data or LLM-generated query-document pair data to perform contrastive training on top of LLMs. LLM2Vec BehnamGhader et al. (2024), on the other hand, conducts further masked next token prediction pre-training with bidirectional attention, and SimCSE Gao et al. (2021) trains on raw text data to transform LLMs into text encoders. Although no labeled data is used, these methods require synthetic or unsupervised paired text data to perform contrastive pre-training (thus still experiencing training costs in terms of computations; and further computational costs may be associated with the generation of synthetic training data). Our method instead relies solely on prompt engineering to transform LLM into a robust text encoder for document retrieval without any extra training.
2.3 Prompting LLMs for document ranking
Inspired by the prompt-following capacity of LLMs, recent studies have explored prompting LLMs for document re-ranking. For instance, UPR Sachan et al. (2022) ranks documents pointwise by prompting the LLM to generate a relevant query for a given document and rank documents based on the likelihood of generating the query. RankGPT Sun et al. (2023) and LRL Ma et al. (2023b) propose to re-rank a list of documents at once and generate permutations for the reordered list. Pairwise Qin et al. (2024) and Setwise Zhuang et al. (2023b) prompting methods have also been explored to improve effectiveness and efficiency in the LLM re-ranking pipeline. These methods are only feasible for re-ranking a handful of candidate documents, thus limited to second-stage document re-ranking. In contrast, our approach utilizes prompts to construct the first-stage retrievers.
@@ 2.4 Prompting LLM for sentence embeddings
The methods most similar to ours prompt LLMs to generate sentence embeddings for semantic textual similarity (STS) tasks Jiang et al. (2023b); Lei et al. (2024); Zhang et al. (2024). These previous methods also used an Explicit One-word Limitation (EOL) prompt, which also instructs LLMs to represent a sentence with one word. However, these methods only evaluate such prompts on STS datasets, and their effectiveness on information retrieval datasets with large document corpora is unknown. Additionally, these methods only represent text with dense embeddings from the hidden states; our method instead generates dense and sparse representations simultaneously to build a hybrid retrieval system. Our empirical results show that dense embeddings alone perform poorly for document retrieval tasks with some LLMs, but sparse representations are much more robust, and the best retrieval effectiveness is achieved with the hybrid retrieval system with scaled model size.
@@ 3 PromptReps
Previous work that leverages LLMs for document ranking are limited to document re-ranking tasks with prompts or rely on contrastive learning to transform a generative LLM into an embedding model for document retrieval. Unlike these previous works, here we aim to directly prompt LLMs to generate both dense embedding representations and sparse bag-of-words representations for document retrieval without any form of extra training effort. To achieve this, we devise the prompt as illustrated in Figure 1 as the input text for LLMs, where <System> <User> and <Assistant> are LLM pre-defined conversational prefix tokens and [text] is the placeholder for passage text.
When using this prompt for text generation, the language model needs to find a single word in its token vocabulary that can best represent the given passage to generate. However, since there could be multiple words to represent the passage, there might be multiple tokens in the vocabulary that have a high probability of being sampled by the language model. Such a distribution over the vocabulary, which is often refers to as “logits”, could provide a good representation of the given passage. In addition, since the logits are computed by the last layer hidden state [^4] Often through dot product between the last hidden state with all token embeddings. of the last input token (‘ ’), which is a dense vector embedding, it could also serve as a dense representation of the passage.
@@
Specifically, during the document indexing phase, we pass all the documents (one at the time) with our prompt into the LLM to get output hidden states and logits. To build a sparse retrieval pipeline with logits, we first need to sparsify the logits representation to be able to perform efficient sparse retrieval. This is because logits originally had values for all tokens in the vocabulary, essentially forming dense vectors with dimensions equal to the vocabulary size. To sparsify the logit representations for sparse retrieval, we perform the following steps:
Table 1: nDCG@10 scores of BEIR 13 publicly available datasets. The best scores of methods without interpolating with BM25 are highlighted in bold.
Unsup Contrastive training PromptReps (ours)
LLM - BERT-330M Llama3-8B-I Llama3-8B-I Llama3-70B-I
Dataset BM25 E5-PTlarge LLM2Vec Dense Sparse Hybrid Dense Sparse Hybrid +BM25
arguana 39.70 44.4 51.73 29.70 22.85 32.98 31.65 24.66 35.27 39.53
climatefever 16.51 15.7 23.58 19.92 9.98 21.38 19.95 12.14 22.18 23.34
dbpedia 31.80 37.1 26.78 31.53 28.84 37.71 31.12 28.30 37.59 41.63
fever 65.13 68.6 53.42 56.28 52.35 71.11 42.06 51.75 63.97 74.06
fiqa 23.61 43.2 28.56 27.11 20.33 32.40 30.80 22.16 34.66 35.35
hotpotqa 63.30 52.2 52.37 19.64 44.75 47.05 24.32 42.12 48.51 65.29
nfcorpus 32.18 33.7 26.28 29.56 28.18 32.98 33.84 29.74 36.08 37.64
nq 30.55 41.7 37.65 34.43 29.55 43.14 38.25 30.37 46.97 48.30
quora 78.86 86.1 84.64 72.55 68.27 80.45 76.14 68.77 82.56 85.83
scidocs 14.90 21.8 10.39 18.51 11.57 17.59 20.59 13.25 19.10 18.82
scifact 67.89 72.3 66.36 52.68 58.48 65.71 63.12 61.53 70.34 73.58
trec-covid 59.47 61.8 63.34 59.52 54.59 69.25 67.64 63.00 76.85 80.29
touche 44.22 19.8 12.82 14.85 18.47 21.78 15.56 18.65 22.35 34.15
avg 43.70 44.61 41.38 35.87 34.48 44.13 38.08 35.88 45.88 50.60
@@
    Finally, the logits are quantized by multiplying the original values by 100 and taking the integer operation on that, and the obtained values represent the weights of corresponding tokens.
With these adjustments, the logits representations of documents are heavily sparsified, allowing for efficient sparse retrieval with an inverted index.
For dense retrieval, we directly use the hidden states as the embeddings of the documents. For indexing these embeddings, we simply normalize all the embeddings and add them into an Approximate Nearest search (ANN) vector index.
At query time, we process the queries exactly the same as the documents, with the only exception being that the term "passage" in the prompt is replaced with "query". The dense representation of the query is utilized for semantic search via the ANN index, while the sparse representation of the query is employed for exact term matching via the inverted index. Following previous work Wang et al. (2021b), we compute the final document scores by applying min-max normalization to both dense and sparse document scores. These normalized scores are then linearly interpolated with equal weights to produce the final document scores.